JoelTheBoss (joel_the_boss):

Find the equation of the quadratic function that passes through the points (-2,18), (3,-2) and (1,0)

JoelTheBoss (joel_the_boss):

Help :/

OpenStudy (zzr0ck3r):

ok so we have \(ax^2+bx+c=0\). use the root \(0=a*1^2+b*1+c=\\a+b+c=0\) then use another point \(18=4a-2b+c\) use the last point \(-2=9a+3b+c\) can you solve this system of three equations to find \(a,c\) and \(c\)?

OpenStudy (zzr0ck3r):

erase the =0 from the first line.

OpenStudy (zzr0ck3r):

and the obvious typo in the last line:)

JoelTheBoss (joel_the_boss):

Thanks a lot :)